Express using a positive exponent.
s^-6
What does it equal?

Answer by wilft1(217)


s^-6 = 1/s^6

likewise, if you had 2^-3 it would become 1/2^3 = 1/2x2x2 = 1/8
you cant multiply s 6 times, so your answer is simply 1/s^6
thats in positive exponent